Bill is 32 years younger than Alicia. 9 years ago, Alicia's age was 3 time bills age . How old is bill ?
Answer by josgarithmetic(39618)

Ages are B and A.
-
Bill is 32 years younger than Alicia, B=32+A.
9 years ago, Alicia's age was 3 time bills age, A-9=3(B-9).
-
Treat those two equations as a system and solve for A and B.